# compat=1.1.0

### v1

@@ Abstract
In warm inflation (WI), the persistent thermal bath that is sustained by dissipative interactions with the inflaton field produces a stochastic background of gravitational waves (GWs). In this paper we study the production and evolution of these GWs. Specifically, we investigate the emission of thermal gravitons (gravitons emitted by a thermal bath) from particle scattering in the bath and the evolution of the corresponding high-energy GWs. We find that the bulk of thermal graviton production in WI occurs during the transition to radiation domination after inflation. Further, the energy density of thermal gravitons is enhanced by roughly one to two orders of magnitude compared to that in a radiation-dominated scenario with the same reheating temperature. We also calculate the spectrum of the resulting stochastic GW background and find that it has a distinctive shape, consisting of a peak at high frequencies ($\sim 100\,{\rm GHz}$) and an almost flat spectrum extending to low frequencies. The peak arises from emission of sub-horizon modes that follow the temperature of the bath. The flat part of the spectrum corresponds to the modes that exit the horizon during WI and re-enter during radiation domination and reflects the approximately constant temperature of the thermal bath during WI. We show that the detection prospects for the high-frequency peak of the GW spectrum, while improved slightly compared to the radiation-dominated case, still remain challenging. The thermal spectrum’s low-frequency plateau is typically subdominant to the amplitude of the standard vacuum tensor modes from inflation, although WI models can exist where the thermal graviton plateau surpasses the vacuum contribution without exceeding current observational limits on the tensor-to-scalar ratio. Furthermore, we calculate the thermal graviton contribution from WI to dark radiation and show that WI models are generally expected to satisfy current observational bounds, including those from the cosmic microwave background (CMB).
@@ I Introduction
The fundamental mechanism underlying this graviton production involves the constituent particles of the system experiencing acceleration due to collisions and scattering off one another, with the rate of scattering depending on their number density and interactions. As a result, the system loses energy through the emission of GWs [4]. Under ordinary conditions in the current Universe, even a fully thermalized system emits only a small amount of energy into gravitons. For reference, the solar gravitational radiation power is $\sim 10^{8}\,\mathrm{W}$ [5, 6], far below the $\sim 10^{26}$ W of electromagnetic radiation power [7]. In the early Universe, however, when the primordial thermal bath could reach much higher temperatures, this emission can be much larger and manifest as a stochastic GW background. There are two regimes for this emission process. In the first regime, the case relevant to this paper, there is sufficient time between collisions to prevent interference, and each collision contributes independently to graviton production. This leads to the microscopic emission of high-energy (or equivalently, high-frequency) gravitons, with typical momenta $k\sim T$, where $T$ is the temperature of the thermal bath. In the second regime, relevant for low-energy graviton emission with $k\ll T$ (not the case in this paper), the system behaves collectively like a fluid and graviton production arises from macroscopic hydrodynamic fluctuations  [4, 8, 9].
@@
Extensive studies have explored the irreducible stochastic GW background sourced by the thermal bath in the radiation-dominated era, assuming the particle content consists solely of the Standard Model (SM) species [8, 9, 11]. Since thermal emission peaks at energies comparable to the temperature, and the frequency of the emitted GWs redshifts in the same way as the temperature, the present-day peak of the GW spectrum is expected to lie in the microwave range; more precisely, around 80 GHz. While the peak always occurs at the same frequency, its amplitude is sensitive to the highest temperature of the bath ever attained (e.g. the reheating temperature). In principle, the resultant stochastic GW background contributes into dark radiation and can affect the expansion rate of the Universe. Observational constraints on dark radiation arise from the predicted abundances of light elements during Big Bang Nucleosynthesis (BBN), and from the decoupling of the cosmic microwave background (CMB) of photons, which occur when the temperature of the Universe is $\sim 0.1\,\text{MeV}$ and $\sim 0.1\,\text{eV}$, respectively. For reasonable choices of the maximum temperature of the SM thermal bath, the contribution from thermal gravitons remains well below these bounds [11]. The effects of additional degrees of freedom motivated by physics beyond the SM [11], hidden sectors, and non-standard cosmologies [12], have also been explored in this context. For instance, it is shown that, in a radiation-dominated epoch, adding new, weakly-coupled degrees of freedom to the SM does not enhance the resulting GW signal; while these particles increase the emission of thermal gravitons, they also contribute to the Hubble expansion rate, leading to greater redshifting of the signal [11, 12].
@@
Here we explore the emission, evolution, and spectrum of GWs from the thermal bath produced during WI. We note that the contribution of hydrodynamic fluctuations to GW emission in the WI context has been previously explored in Refs. [13, 14]. During the WI inflationary period, thermal gravitons are continually produced but then redshifted away due to the quasi-exponential expansion of the Universe. Indeed we will show that the production of thermal gravitons by WI peaks during the transition to the radiation domination phase and their energy density is roughly one to two orders of magnitude larger than that of the GWs emitted during radiation domination with the same reheating temperature. By calculating the spectrum of the associated stochastic GW background, we will show that it has a distinctive shape which consists of a sharp peak at high frequencies, followed by an almost flat part extending to low frequencies. The peak represents the emission of sub-horizon modes that follow the temperature of the bath, while the flat part originates from the modes that exit the horizon during WI and re-enter during radiation domination. Furthermore, we explore detection prospects of the thermal stochastic GW background produced by WI and find that the high-frequency peak, while enhanced compared to the radiation-dominated case, still remains challenging to detect. We also investigate the conditions on the evolution of the temperature and the Hubble expansion rate during WI that enable the low-frequency plateau of the thermal spectrum to surpass the vacuum tensor spectrum, while keeping the tensor-to-scalar ratio within current observational limits. Finally, we determine the conditions under which observations can constrain the contribution to dark radiation of thermal gravitons produced by WI.
@@ II Warm Inflation: A Brief Review
where $\rho_{\phi}=V(\phi)+\dot{\phi}^{2}/2$ is the inflaton energy density, $\rho_{r}=(\pi^{2}/30)g_{\star}(T)T^{4}$ is the energy density of the thermal bath with $g_{\star}(T)$ number of relativistic degrees of freedom, and $M_{\rm pl}\equiv 1/\sqrt{8\pi G}\approx 2.4\times 10^{18}\,$eV. The Hubble expansion rate $H$ is determined by the Friedmann equation. The dissipation term, $\Upsilon(\phi,T)$, characterizes the rate at which the inflaton field’s energy is transferred into radiation and its specific form is determined by the details of the underlying microphysical model [3].
For the remainder of this work, we aim to keep our discussion general and model-independent. However, to support our results with a concrete example, we consider a quartic inflaton potential, $V(\phi)=\lambda\phi^{4}$ along with a linear dissipation rate $\Upsilon\propto T$. This choice is motivated by the Warm Little Inflaton model [15, 16], which provides the simplest successful realization of WI consistent with CMB observations. It is customary to introduce the dimensionless dissipation strength $Q$, which characterizes the efficiency with which the inflaton transfers its energy into radiation:
@@ III.1 Energy Density of Gravitational Waves
$\frac{d\rho_{\rm GW}}{dt}+4H\rho_{\rm GW}=\frac{4T^{4}}{M^{2}_{\rm Pl}}\int\frac{d^{3}k}{(2\pi)^{3}}\hat{\eta}(T,\hat{k}),$ (5)
where the dimensionless source term, $\hat{\eta}(T,\hat{k})$, encodes the production rate and kinematic distribution of gravitons from particle scatterings in the thermal plasma at a temperature $T$, and $\hat{k}\equiv k/T$ where $k$ is the physical momentum. The exact form of $\hat{\eta}(T,\hat{k})$ depends on the underlying microphysics, including the particle content and interaction rates of the thermal bath, and has been studied for a generic quantum field theory coupled to gravity in Refs. [8, 9, 11]. The leading-log estimate of $\hat{\eta}(T,\hat{k})$ around its peak can be expressed as [8, 11]:
@@
where $A_{\eta}(T)$ captures the dependence on the Debye thermal masses of the bath particles, normalized by the temperature, and is expected to be $\sim\mathcal{O}(1)$. A full leading order calculation of $\hat{\eta}(T,\hat{k})$ can be found in Ref. [9], which includes logarithmic terms that depend on $\hat{k}$. However, without loss of generality, and to maintain a model-independent set-up for our calculations, we use the leading-log estimate of $\hat{\eta}(T,\hat{k})$, i.e., Eq. (6), with $A_{\eta}(T_{\rm max})=1$ for the rest of this work. It is worth noting that, as expected, the integrand that governs the graviton production, $\hat{k}^{2}\hat{\eta}(T,\hat{k})$, peaks at a physical momentum of the order of the thermal bath temperature, specifically for $k\simeq 2.82\,T$.
@@ III.2 Spectrum of Gravitational Waves
×$\times\int_{N_{e,{\rm ini}}}^{N_{e}}\frac{d\mathcal{I}_{\rm GW}}{d\ln k}[k(N_{e}^{\prime}),N_{e}^{\prime}]\,dN_{e}^{\prime},$ (15)
@@
×$\times\Omega_{\gamma,0}\,\Omega_{\rm GW}[k(N^{*}_{e}),N^{*}_{e}],$ (17)
@@
×$\times\left(\frac{g_{\star,s}(T_{0})}{g_{\star,s}[T(N^{\rm re}_{e})]}\right)^{\frac{4}{3}}\Omega_{\gamma,0}\,\mathcal{P}^{\rm th}_{T}(k_{c}),$ (22)
@@
Here we used the fact that a WI setting must satisfy the condition $H<T$ throughout. Hence at horizon crossing, where $k=H$, $\hat{k}=H/T\ll 1$ so that we may approximate $[\hat{k}^{3}\hat{\eta}](N_{e}^{\rm ex})\approx\hat{k}^{3}(N_{e}^{\rm ex})=[H(N_{e}^{\rm ex})/T(N_{e}^{\rm ex})]^{3}$.
@@
The high-frequency peak arises from the sub-horizon modes. Such peak is expected to be higher than the corresponding thermal graviton spectrum from a radiation-dominated era with the same reheating temperature, reflecting the same enhancement factor introduced in the context of the total GW energy density, i.e. $R_{\rm GW}$ in Eq. (13). There is also a mild peak-frequency shift relative to the GW spectrum from a radiation-dominated era due to the factor $[T(N_{e}^{*})e^{N_{e}^{*}}]/[T(N_{e}^{\Omega})e^{N_{e}^{\Omega}}]$ in Eq. (19). In fact, while this term equals unity in a radiation-dominated Universe, in a WI setting the non-zero entropy injection prior to radiation domination makes it slightly larger than one, shifting the peak of the GW spectrum to lower frequencies. However, as we will show below, this effect is negligible in practice.
@@ IV.1 Constraints and Detection Prospects for the thermal GW spectrum
The enhanced high-frequency peak presents an interesting target for detection efforts, though with challenging prospects. In recent years, searches for high-frequency GWs above the kHz range have attracted vast interest (see [26] for a review of recent developments). Search strategies typically rely on the inverse Gertsenshtein effect [27], where an effective conversion of GWs into electromagnetic fields occurs as they propagate through an external magnetic field. However, the probability of graviton conversion is suppressed by $M^{2}_{\rm pl}$. Nevertheless exclusion limits have been set based on the non-detection of conversion in the magnetic fields in astrophysical environments [28, 29, 30, 31, 32, 33, 34, 35, 36, 37], as well as in existing and proposed laboratory facilities [38, 39, 40, 41, 42, 43, 44, 45, 46]. Ref. [11] has shown that, even under optimistic projections using the Gaussian beam technique in future graviton/photon conversion experiments, a reheating temperature near the Planck scale is required in the case of instantaneous reheating into a radiation-dominated universe. In the context of thermal graviton detection, the thermal GW signal we derived here within a WI setting only offers a modest improvement over this challenging baseline, with an enhancement factor $R_{\rm GW}\sim 30$ relative to instantaneous reheating scenarios to a radiation-dominated era with the same initial temperature. Consequently, the detection prospects of this high-frequency signal ultimately depend on the viability to achieve a sufficiently high reheating temperature within a realistic WI framework. Below we describe the bounds set by $\Delta N_{\rm eff}$ measurements on the graviton energy density, which are far more constraining than any existing bounds from graviton/photon conversion.
@@
×$\times\,\frac{1}{n\left(n+4\,\Delta N_{e}^{\rm ex}\right)}\frac{Q}{1+Q},$ (26)
[…]
Figure 3: Ratio of the low-frequency thermal and vacuum GW spectra, $\Omega_{\rm GW,0}^{\rm th}/\Omega_{\rm GW,0}^{\rm vac}$ as a function of the Hubble parameter $H$ and temperature $T$ during inflation. The color scale to the right indicates the magnitude of $\Omega_{\rm GW,0}^{\rm th}/\Omega_{\rm GW,0}^{\rm vac}$. The dot-dashed green line in the figure marks the boundary at which the thermal contributions equal those from vacuum tensor perturbations; in the yellow region above the line, the thermal contribution dominates. The brown shaded region corresponds to $H>T$, where WI is not viable. Black and gray dashed lines represent current and future constraints on the tensor-to-scalar ratio $r$ at $2\sigma$, respectively from the combination of Planck and BICEP/Keck measurements [47], and CMB-S4 forecasts [48]. Note that these constraints apply specifically to CMB modes, which exit the horizon approximately 60 e-folds before the end of inflation. The region of particular interest lies between the green contour and the constraint lines, where thermal contributions dominate over vacuum ones while remaining consistent with current observational bounds and within reach of future experiments such as CMB-S4. For reference, we indicate as red ($Q_{\rm ini}=10^{-2}$) and blue ($Q_{\rm ini}=1$) diamonds the corresponding values of $\Omega_{\rm GW,0}^{\rm th}/\Omega_{\rm GW,0}^{\rm vac}$ for the two representative WI examples analyzed in this work. As a whole, the figure highlights the fundamental challenge of observationally distinguishing WI signatures through thermal GWs in realistic scenarios, requiring “fine-tuned” combinations of $H,\,T$ to achieve both thermal dominance and observational accessibility.

### v2

# Thermal Gravitons from Warm Inflation

@@ Abstract
In warm inflation (WI), the persistent thermal bath that is sustained by dissipative interactions with the inflaton field produces a stochastic background of gravitational waves (GWs). In this paper we study the production and evolution of these GWs. Specifically, we investigate the emission of thermal gravitons (gravitons emitted by a thermal bath) from scattering with particles in the bath and then the evolution of the corresponding high-frequency GWs. We find that the bulk of thermal graviton production in WI occurs during the transition to radiation domination after inflation. Further, the energy density of thermal gravitons is enhanced by roughly one to two orders of magnitude compared to that in a radiation-dominated scenario with the same reheating temperature. We also calculate the spectrum of the resulting stochastic GW background and find that it has a distinctive shape, consisting of a peak at high frequencies ($\sim 100\,{\rm GHz}$) and an almost flat spectrum extending to low frequencies. The peak arises from emission of sub-horizon modes that follow the temperature of the bath. The flat part of the spectrum corresponds to the modes that exit the horizon during WI and re-enter during radiation domination and reflects the approximately constant temperature of the thermal bath during WI. We show that the detection prospects for the high-frequency peak of the GW spectrum, while improved slightly compared to the radiation-dominated case, still remain challenging. The thermal spectrum’s low-frequency plateau is typically subdominant to the amplitude of the standard vacuum tensor modes from inflation, although WI models can exist where the thermal graviton plateau surpasses the vacuum contribution without exceeding current observational limits on the tensor-to-scalar ratio. Furthermore, we calculate the thermal graviton contribution from WI to dark radiation and show that WI models are generally expected to satisfy current observational bounds including those from the cosmic microwave background (CMB).
@@ I Introduction
The fundamental mechanism underlying this graviton production involves the constituent particles of the system experiencing acceleration due to collisions and scattering off one another, with the rate of scattering depending on their number density and interactions. As a result, the system loses energy through the emission of GWs [4]. Under ordinary conditions in the current Universe, even a fully thermalized system emits only a small amount of energy into gravitons. For reference, the solar gravitational radiation power is $\sim 10^{8}\,\mathrm{W}$ [5, 6], far below the $\sim 10^{26}$ W of electromagnetic radiation power [7]. In the early Universe, however, when the primordial thermal bath could reach much higher temperatures, this emission can be much larger and manifest as a stochastic GW background. There are two regimes for this emission process. In the first regime, the case relevant to this paper, there is sufficient time between collisions to prevent interference, and each collision contributes independently to graviton production. This leads to the microscopic emission of high-energy (or equivalently, high-frequency) gravitons, with typical momenta $k\sim T$, where $T$ is the temperature of the thermal bath. In the second regime, relevant for low-energy graviton emission with $k\ll T$ (not the case in this paper), the system behaves collectively like a fluid. In this regime, graviton production arises from macroscopic hydrodynamic fluctuations  [4, 8, 9], though this contribution is negligible in the WI context and will not be considered further in this work.
@@
Extensive studies have explored the irreducible stochastic GW background sourced by the thermal bath in the radiation-dominated era, assuming the particle content consists solely of the Standard Model (SM) species [8, 9, 11]. Since thermal emission peaks at energies comparable to the temperature, and the frequency of the emitted GWs redshifts in the same way as the temperature, the present-day peak of the GW spectrum is expected to lie in the microwave range; more precisely, around 80 GHz. While the peak always occurs at the same frequency, its amplitude is sensitive to the highest temperature of the bath ever attained (e.g. the reheating temperature). In principle, the resultant stochastic GW background contributes into dark radiation and can affect the expansion rate of the Universe. Observational constraints on dark radiation arise from the predicted abundances of light elements during Big Bang Nucleosynthesis (BBN), and from the decoupling of the cosmic microwave background (CMB) of photons, which occur when the temperature of the Universe is $\sim 0.1\,\text{MeV}$ and $\sim 0.1\,\text{eV}$, respectively. For reasonable choices of the maximum temperature of the SM thermal bath, the contribution from thermal gravitons remains well below these bounds [11]. The effects of additional degrees of freedom motivated by physics beyond the SM [11], hidden sectors [12], and non-standard cosmologies [13, 14, 15], have also been explored in this context. For instance, it is shown that, in a radiation-dominated epoch, adding new, weakly-coupled degrees of freedom to the SM does not enhance the resulting GW signal; while these particles increase the emission of thermal gravitons, they also contribute to the Hubble expansion rate, leading to greater redshifting of the signal [11, 13].
@@
Here we explore the emission, evolution, and spectrum of GWs from microscopic particle collisions in the thermal bath produced during WI. We note that the contribution of hydrodynamic fluctuations to GW emission in the WI context has been previously explored in Refs. [16, 17], where it was shown to exhibit the universal $\sim k^{3}$ spectral shape with a model-dependent magnitude controlled by the shear viscosity. During the WI inflationary period, thermal gravitons are continually produced but then redshifted away due to the quasi-exponential expansion of the Universe. Indeed we will show that the production of thermal gravitons by WI peaks during the transition to the radiation domination phase and their energy density is roughly one to two orders of magnitude larger than that of the GWs emitted during radiation domination with the same reheating temperature. By calculating the spectrum of the associated stochastic GW background, we will show that it has a distinctive shape which consists of a sharp peak at high frequencies, followed by an almost flat part extending to low frequencies. The peak represents the emission of sub-horizon modes that follow the temperature of the bath, while the flat part originates from the modes that exit the horizon during WI and re-enter during radiation domination. Furthermore, we explore detection prospects of the thermal stochastic GW background produced by WI and find that the high-frequency peak, while enhanced compared to the radiation-dominated case, still remains challenging to detect. We also investigate the conditions on the evolution of the temperature and the Hubble expansion rate during WI that enable the low-frequency plateau of the thermal spectrum to surpass the vacuum tensor spectrum, while keeping the tensor-to-scalar ratio within current observational limits. Finally, we determine the conditions under which observations can constrain the contribution to dark radiation of thermal gravitons produced by WI.
@@ II Warm Inflation: A Brief Review
where $\rho_{\phi}=V(\phi)+\dot{\phi}^{2}/2$ is the inflaton energy density, $\rho_{r}=(\pi^{2}/30)g_{\star}(T)T^{4}$ is the energy density of the thermal bath with $g_{\star}(T)$ number of relativistic degrees of freedom, and $M_{\rm pl}\equiv 1/\sqrt{8\pi G}\approx 2.4\times 10^{18}\,$GeV. The Hubble expansion rate $H$ is determined by the Friedmann equation. The dissipation term, $\Upsilon(\phi,T)$, characterizes the rate at which the inflaton field’s energy is transferred into radiation and its specific form is determined by the details of the underlying microphysical model [3].
For the remainder of this work, we aim to keep our discussion general and model-independent. However, to support our results with a concrete example, we consider a quartic inflaton potential, $V(\phi)=\lambda\phi^{4}$ along with a linear dissipation rate $\Upsilon\propto T$. This choice is motivated by the Warm Little Inflaton model [18, 19], which provides the simplest successful realization of WI consistent with CMB observations, as well as by recent developments involving an axion-like inflaton field coupled solely through Standard Model gauge interactions [20]. We stress, however, that the key qualitative features of our findings remain independent of any specific model choice.
It is customary to introduce the dimensionless dissipation strength $Q$, which characterizes the efficiency with which the inflaton transfers its energy into radiation:
@@ III.1 Energy Density of Gravitational Waves
[…]
@@
[…]
@@ III.2 Spectrum of Gravitational Waves
$\times\int_{N_{e,{\rm ini}}}^{N_{e}}\frac{d\mathcal{I}_{\rm GW}}{d\ln k}[k(N_{e}^{\prime}),N_{e}^{\prime}]\,dN_{e}^{\prime},$ (15)
@@
$\times\Omega_{\gamma,0}\,\Omega_{\rm GW}[k(N^{*}_{e}),N^{*}_{e}],$ (17)
@@
$\times\left(\frac{g_{\star,s}(T_{0})}{g_{\star,s}[T(N^{\rm re}_{e})]}\right)^{\frac{4}{3}}\Omega_{\gamma,0}\,\mathcal{P}^{\rm th}_{T}(k_{c}),$ (22)
@@
Here we used the fact that a WI setting must satisfy the condition $H<T$ throughout. Hence at horizon crossing, where $k=H$, $\hat{k}=H/T\ll 1$ so that we may approximate $\hat{k}^{3}\hat{\eta}](N_{e}^{\rm ex})\approx\hat{k}^{3}(N_{e}^{\rm ex})=[H(N_{e}^{\rm ex})/T(N_{e}^{\rm ex})]^{3}$.
@@
The high-frequency peak arises from the sub-horizon modes. Such peak is expected to be higher than the corresponding thermal graviton spectrum from a radiation-dominated era with the same reheating temperature, reflecting the same enhancement factor introduced in the context of the total GW energy density, i.e. $R_{\rm GW}$ in Eq. (13). There is also a mild peak-frequency shift relative to the GW spectrum from a radiation-dominated era due to the factor $T(N_{e}^{*})e^{N_{e}^{*}}]/[T(N_{e}^{\Omega})e^{N_{e}^{\Omega}}]$ in Eq. (19). In fact, while this term equals unity in a radiation-dominated Universe, in a WI setting the non-zero entropy injection prior to radiation domination makes it slightly larger than one, shifting the peak of the GW spectrum to lower frequencies. However, as we will show below, this effect is negligible in practice.
@@ IV.1 Constraints and Detection Prospects for the thermal GW spectrum
Figure 3: Ratio of the low-frequency thermal and vacuum GW spectra, $\Omega_{\rm GW,0}^{\rm th}/\Omega_{\rm GW,0}^{\rm vac}$ as a function of the Hubble parameter $H$ and temperature $T$ during inflation. The color scale to the right indicates the magnitude of $\Omega_{\rm GW,0}^{\rm th}/\Omega_{\rm GW,0}^{\rm vac}$. The dot-dashed green line in the figure marks the boundary at which the thermal contributions equal those from vacuum tensor perturbations; in the yellow region above the line, the thermal contribution dominates. The brown shaded region corresponds to $H>T$, where WI is not viable. Black and gray dashed lines represent current and future constraints on the tensor-to-scalar ratio $r$ at $2\sigma$, respectively from the combination of Planck and BICEP/Keck measurements [30], and CMB-S4 forecasts [31]. Note that these constraints apply specifically to CMB modes, which exit the horizon approximately 60 e-folds before the end of inflation. The region of particular interest lies between the green contour and the constraint lines, where thermal contributions dominate over vacuum ones while remaining consistent with current observational bounds and within reach of future experiments such as CMB-S4. For reference, we indicate as red ($Q_{\rm ini}=10^{-2}$) and blue ($Q_{\rm ini}=1$) diamonds the corresponding values of $\Omega_{\rm GW,0}^{\rm th}/\Omega_{\rm GW,0}^{\rm vac}$ for the two representative WI examples analyzed in this work. As a whole, the figure highlights the fundamental challenge of observationally distinguishing WI signatures through thermal GWs in realistic scenarios, requiring “fine-tuned” combinations of H,TH,\,T to achieve both thermal dominance and observational accessibility.
@@
The enhanced high-frequency peak presents an interesting target for detection efforts, though with challenging prospects. In recent years, searches for high-frequency GWs above the kHz range have attracted vast interest (see [32] for a review of recent developments). Search strategies typically rely on the inverse Gertsenshtein effect [33], where an effective conversion of GWs into electromagnetic fields occurs as they propagate through an external magnetic field. However, the probability of graviton conversion is suppressed by $M^{2}_{\rm pl}$. Nevertheless exclusion limits have been set based on the non-detection of conversion in the magnetic fields in astrophysical environments [34, 35, 36, 37, 38, 39, 40, 41, 42, 43], as well as in existing and proposed laboratory facilities [44, 45, 46, 47, 48, 49, 50, 51, 52]. Ref. [11] has shown that, even under optimistic projections using the Gaussian beam technique in future graviton/photon conversion experiments, a reheating temperature near the Planck scale is required in the case of instantaneous reheating into a radiation-dominated universe. In the context of thermal graviton detection, the thermal GW signal we derived here within a WI setting only offers a modest improvement over this challenging baseline, with an enhancement factor $R_{\rm GW}\sim 30$ relative to instantaneous reheating scenarios to a radiation-dominated era with the same initial temperature. Consequently, the detection prospects of this high-frequency signal ultimately depend on the viability to achieve a sufficiently high reheating temperature within a realistic WI framework. Given the well-defined peak frequency of the thermal GW spectrum, a complementary strategy is to employ dedicated resonant searches targeting the signal around $10^{11}\,{\rm Hz}$. For instance, Ref. [53] explores a resonant cavity design capable of reaching a characteristic strain sensitivity of $\mathcal{O}(10^{-36})$ at $10^{8}\,{\rm Hz}$. With further optimization toward the THz band, such resonant setups could provide a promising avenue to probe WI thermal gravitons. Below we describe the bounds set by $\Delta N_{\rm eff}$ measurements on the graviton energy density, which are far more constraining than any existing bounds from graviton/photon conversion.
@@
$\times\,\frac{1}{n\left(n+4\,\Delta N_{e}^{\rm ex}\right)}\frac{Q}{1+Q},$ (26)
where $\Delta N_{e}^{\rm ex}\equiv(N^{\rm infl}_{e}-N_{e}^{\rm ex})>0$, with $N_{e}^{\rm infl}=60$. [^4] Here $N_{e}^{\rm infl}=60$ denotes the end of inflation, and we set initial conditions at $N_{e}^{\rm ex}=0$, corresponding approximately to the horizon exit of CMB modes. Note that this choice of e-fold counting is arbitrary and does not affect physical results. The factor in the second line is evidently always less than unity for $n\gtrsim\mathcal{O}(1)$, thus ensuring that the thermal contributions to the primordial spectrum remain highly suppressed in minimal WI models.